### v7

# Exploring Cross-model Neuronal Correlations in the Context of Predicting Model Performance and Generalizability

Abstract
As Artificial Intelligence (AI) models are increasingly integrated into critical systems, the need for a robust framework to establish the trustworthiness of AI is increasingly paramount. While collaborative efforts have established conceptual foundations for such a framework, there remains a significant gap in developing concrete, technically robust methods for assessing AI model quality and performance. This paper introduces a novel approach for assessing a newly trained model’s performance based on another known model by calculating correlation between neural networks. The proposed method evaluates correlations by determining if, for each neuron in one network, there exists a neuron in the other network that produces similar output. This approach has implications for memory efficiency, allowing for the use of smaller networks when high correlation exists between networks of different sizes. Experiments on five fully connected networks and a two layer CNN trained on MNIST family datasets show that higher alignment with the CNN tracks stronger performance and smaller degradation under black box transfer based attacks. On ImageNet pretrained ResNets and DenseNets, partial layer comparisons recover intuitive architectural affinities, indicating that the procedure scales with reasonable approximations. These results support representational alignment as a lightweight compatibility check that complements standard accuracy, calibration, and robustness evaluations and enables early external validation of new models. Code is available at https://github.com/aheldis/Cross-model-Correlation.git.
keywords:
Trustworthy AI; external validation; training data independent evaluation; representational alignment; neuronal correlation
1 Introduction
Artificial intelligence (AI) has moved from supporting roles to the core of what many societies regard as essential services, from healthcare to public safety [guilherme2019ai]. As these systems influence increasingly consequential decisions, their validation can no longer rely on ad hoc checks or narrow task accuracy alone. Robust, transparent approaches are needed to establish whether a model behaves reliably under real-world conditions and evolving inputs [batarseh2021survey, sujan2023validation]. In parallel, regulators and licensing bodies are signaling that AI systems will need oversight and audit practices commensurate with other high-stakes technologies [clark2019regulatory, o2019legal].
Today’s validation toolkits remain heavily dependent on developer-controlled ingredients—training and validation data, simulation harnesses, and expert judgments—which are invaluable yet inherently internal [myllyaho2021systematic]. Post-market monitoring adds an important safety net, but by definition it observes failures only after deployment [myllyaho2021systematic]. What is missing is a complementary pathway for external and independent assessment that does not require privileged access to training data, internal agents, or proprietary evaluation suites.
Motivated by this gap, we explore whether representational alignment between a candidate model and a well-audited reference model can serve as a practical, computable indicator of trustworthiness. Intuitively, if two networks encode inputs similarly across their layers, they may exhibit related strengths and failure modes; conversely, stark representational divergence can be an early warning that a new model is operating outside familiar behavior regimes. We operationalize this idea with a simple cross-model neuronal correlation score designed to be computed from activations alone.
Concretely, given a small probe dataset used only to elicit activations, we compute, for each neuron in one model, the best-correlated neuron in the other (using absolute correlation), apply a depth-aware penalty to respect architectural hierarchy, and then average scores bidirectionally. The result is a single scalar in $[0,1]$ that increases with representational alignment. For modern architectures, we estimate this efficiently by comparing corresponding layers or stages and, where necessary, subsampling neurons. Crucially, the procedure requires no access to training data or internal training artifacts—supporting independent evaluation alongside existing, developer-centric methods [batarseh2021survey, sujan2023validation, myllyaho2021systematic]. [^1] Throughout, we use “training data independent” to mean that our procedure does not require access to a model’s training data or labels, nor to proprietary evaluation suites. A small, unlabeled probe set is still used solely to elicit activations; therefore the method is not strictly “data free”, and estimates can vary with the probe distribution.
We evaluate this metric in two regimes. First, across five fully connected networks (FCNs) and a two-layer CNN trained on MNIST-family datasets, we find that higher cross-model correlation with the CNN generally accompanies stronger performance under transferable, black-box perturbations, while the least-aligned FCN also exhibits the greatest degradation. Second, on ImageNet pretrained ResNets, DenseNets, and EfficientNets, partial-layer analyses recover intuitive architectural affinities (for example, depth-adjacent variants aligning most), indicating that the score scales to contemporary models with tractable approximations.
Contributions.
  1. 1.
    We propose a simple, symmetric neuronal-correlation metric with a layer-aware penalty that can be computed without access to training data.
  2. 2.
    We show, across multiple MNIST-family settings, that cross-model similarity often tracks robustness under transferable black-box attacks.
  3. 3.
    We demonstrate a tractable partial-correlation procedure on large ImageNet models that recovers plausible architectural relationships, supporting the metric’s utility at scale.
Together, these results position cross-model neuronal correlation as a lightweight compatibility check: a model-agnostic indicator linking representational alignment to empirical robustness, designed to complement existing validation practices and emerging regulatory expectations [clark2019regulatory, o2019legal].
2 Preliminaries and Related Work
2.1 Trustworthy AI
As AI systems increasingly inform consequential decisions, trustworthiness hinges on whether models perform reliably beyond narrow i.i.d. test sets and in the presence of realistic shifts and perturbations. In technical terms, this entails (i) strong task performance and calibrated uncertainty on nominal data, and (ii) robustness under distributional variation and adversarial manipulation. Because developer controlled evaluations (for example, internal test suites) can miss important failure modes, there is growing interest in complementary external checks that do not require privileged access to training data or proprietary tooling.
Performance and reliability.
Task accuracy alone is insufficient in high stakes settings because overconfident errors can be harmful. Modern neural networks often exhibit probability miscalibration, where predicted confidences do not match empirical correctness; post hoc methods such as temperature scaling partly address this but have limitations [guo2017calibration]. Subsequent work refined how calibration is measured and documented metric pathologies that can alter conclusions [nixon2019measuring]. Large scale studies further show that calibration depends on model family and evaluation distribution, motivating reliability metrics (for example, NLL, Brier, ECE) alongside accuracy and across test conditions [minderer2021revisiting].
Robustness under distribution shift.
Deployed models face natural shifts in sensors, environments, populations, and time. Benchmarks such as ImageNet C and ImageNet P quantify degradation under common corruptions and small perturbations [hendrycks2019imagenetc], while large testbeds show that robustness to synthetic perturbations does not transfer to natural distribution shifts [taori2020natural]. Predictive uncertainty is also brittle under shift, with many popular methods degrading in reliability [ovadia2019trust]. Beyond image classification, WILDS benchmark aggregates realistic, domain specific shifts across modalities to standardize evaluation [koh2021wilds].
Connection to this work.
Because accuracy, calibration, and robustness each capture different facets of trustworthiness, and because developer curated tests cannot cover all regimes, we investigate an external, training data independent proxy: representational alignment with a well audited reference model. Prior work compares learned representations via SVCCA, PWCCA, and linear CKA [raghu2017svcca, morcos2018pwcca, kornblith2019cka], as well as Representational Similarity Analysis from neuroscience [kriegeskorte2008rsa]. We complement these with a simple, neuron level correlation score computed from activations on a small, unlabeled probe set, designed to serve as a lightweight indicator alongside standard performance and robustness evaluations.
2.2 Model Robustness
Robustness asks whether a model maintains acceptable performance under plausible perturbations and shifts. Two complementary regimes are commonly studied.
Adversarial robustness.
Worst case perturbations expose vulnerabilities even for high accuracy models. Gradient based attacks such as FGSM, DeepFool, and PGD demonstrate that imperceptible perturbations can flip predictions [goodfellow2014explaining, deepfool, madry2018towards], while stronger evaluations and adaptive attacks highlight the brittleness of many defenses [carlini2017towards]. Of particular relevance to external assessment, adversarial examples transfer between independently trained models, which enables black box attacks [papernot2016transfer, tramer2018ensemble]. Transfer is often attributed to shared, non robust features learned by standard training [ilyas2019adversarial], which implies that internal representations across models are partially aligned.
Natural shift robustness.
Beyond worst case bounds, models encounter real world changes in noise, blur, weather, sensors, or demographics. ImageNet C and ImageNet P evaluate corruption and perturbation robustness [hendrycks2019imagenetc], while broad testbeds show that success on synthetic corruptions does not guarantee robustness to natural shifts [taori2020natural]. Uncertainty estimation degrades under shift, which complicates safe deployment [ovadia2019trust]. WILDS provides diverse, realistic out of distribution settings to benchmark progress [koh2021wilds].
Connection to our algorithm.
The widespread transferability of attacks suggests that representational geometry is a key mediator of robustness. Models with more aligned features tend to share both strengths and failure modes. Our method quantifies such alignment directly through a simple, layer aware neuronal correlation score. We use this score to compare candidate models against well audited references, hypothesizing and empirically exploring that stronger representational alignment correlates with better robustness indicators (for example, lower degradation under black box attacks and common corruptions). This external, activation based metric is designed to complement standard robustness evaluations and to support early stage, training data independent assessments.
2.3 Representational Similarity and External Validation
Representation comparison methods study how similarly different networks encode inputs. Early analyses used canonical correlation based approaches such as SVCCA and PWCCA to compare subspaces across layers and training runs [raghu2017svcca, morcos2018pwcca]. Centered Kernel Alignment provides a simple, scalable similarity measure that correlates with transfer performance across architectures [kornblith2019cka]. In parallel, Representational Similarity Analysis compares representational dissimilarity matrices and has linked deep network features to neural and human similarity judgments [kriegeskorte2008rsa]. These methods primarily assess layer or subspace level similarity.
Our work is complementary and intentionally minimalist. Instead of subspace alignment, we compute a direct, neuron level correlation score with a lightweight depth penalty and symmetric averaging across models. This choice makes the metric easy to compute from activations only, using a small unlabeled probe set, and it enables external, training data independent auditing. Conceptually, this aligns with observations from adversarial transfer that independently trained models often learn similar non robust features [papernot2016transfer, ilyas2019adversarial]. Practically, we use the score as a compatibility check by comparing a candidate model to a well audited reference and studying how alignment relates to accuracy and robustness indicators in controlled experiments.
3 Proposed Algorithm
This section describes the proposed method for quantifying the similarity between two trained neural networks based on their internal representations. The goal is to produce a symmetric, data-independent measure that reflects how similarly two models process and encode inputs.
3.1 Setup and Notation
Let two networks $F$ and $G$ map an input $x\in\mathcal{X}$ to a hierarchy of hidden-layer activations. A probe dataset $\mathcal{D}=\{x_{m}\}_{m=1}^{M}$, typically composed of unlabeled validation or public test inputs, is used solely to elicit these activations. For each neuron $u$ in $F$, its activation vector is defined as $\alpha_{u}=[h_{u}(x_{1}),h_{u}(x_{2}),\ldots,h_{u}(x_{M})],$ where $h_{u}(x_{m})$ denotes the scalar activation of neuron $u$ for input $x_{m}$. The same is done for all neurons in $G$. The similarity between two neurons’ activation vectors is measured by the Pearson correlation coefficient $\rho(\cdot,\cdot)$.
3.2 Per-Neuron Best-Match Score
For every neuron $u$ in $F$, we identify the neuron in $G$ whose activation pattern is most strongly correlated with it. This neuron, denoted $v^{*}(u)$, is found as
$v^{*}(u)=\arg\max_{v\in\mathcal{U}_{G}}|\rho(\alpha_{u},\alpha_{v})|,$
where $\mathcal{U}_{G}$ is the set of all hidden neurons in $G$. The absolute value accounts for possible sign inversions caused by subsequent linear transformations or normalization layers.
To incorporate architectural depth information, we introduce a layer-distance penalty. The function $\text{layer}(\cdot)$ returns the integer index of the layer that a neuron belongs to. In multilayer perceptrons (MLPs), layers are numbered sequentially starting from the first hidden layer. In convolutional networks (CNNs), each convolutional or pooling block can be indexed as a layer, and in hierarchical architectures such as ResNets or DenseNets, each stage or transition block may share a single index to represent comparable levels of abstraction.
The layer-distance-normalized correlation score for neuron $u$ is defined as
$S(u;F\rightarrow G)=\frac{|\rho(\alpha_{u},\alpha_{v^{*}(u)})|}{1+|\mathrm{layer}(u)-\mathrm{layer}(v^{*}(u))|}.$
The denominator penalizes matches between neurons that are far apart in depth, ensuring that correspondences between early- and late-layer neurons contribute less to the final measure. The same computation is performed symmetrically for each neuron $v$ in $G$, identifying its best match in $F$.
3.3 Network-Level Correlation
The overall cross-model neuronal correlation is obtained by averaging per-neuron scores from both directions. This ensures that the result is invariant to which model is considered the reference:
$\mathrm{Corr}(F,G)=\frac{1}{2}\Bigg(\frac{1}{|\mathcal{U}_{F}|}\sum_{u\in\mathcal{U}_{F}}S(u;F\rightarrow G)+\frac{1}{|\mathcal{U}_{G}|}\sum_{v\in\mathcal{U}_{G}}S(v;G\rightarrow F)\Bigg).$
The resulting scalar lies in the interval $[0,1]$. Higher values indicate stronger representational similarity between networks, while lower values suggest that the two models have developed distinct internal structures.
3.4 Partial Correlation for Tractability
The computational complexity of full cross-layer matching is quadratic in the number of neurons, $\mathcal{O}(|\mathcal{U}_{F}||\mathcal{U}_{G}|)$, which is impractical for modern architectures with millions of activations. To make the computation feasible, we use a partial correlation strategy. The comparison is restricted to corresponding or functionally similar layers (for instance, comparing stage 4 of a ResNet with transition block 3 of a DenseNet) rather than all possible cross-layer pairs. Additionally, a random subset of neurons can be sampled from each layer to further reduce computational load.
This approximation substantially reduces cost while preserving the ability to detect meaningful alignment between two architectures’ representational spaces. Empirically, partial correlation values remain stable across reasonable subsampling levels and layer selection schemes.
3.5 Black-Box Adversarial Robustness
Additionally, we explore the vulnerability of neural networks to black-box adversarial attacks and examine how this susceptibility relates to our proposed metric. To investigate the architectural relationships between different neural networks, we adapted the attack method introduced by (blackbox), with slight modifications to their original approach. This black-box method facilitates model comparison without necessitating prior knowledge of their architectures or training datasets. As described in Algorithm 1, at each iteration, we train a simple model like a two-layer CNN that differs from the other models used in the experiments on the dataset. Subsequently, we adjust the data using the perturbation computed via the Fast Gradient Sign Method (FGSM) fgsm. We iterate through the process multiple times to generate our ultimate adversaries. In our method, we solely utilize test data instead of custom-crafted data, as the former remains untouched during training. Furthermore, we refrain from altering labels based on the oracle’s feedback and discard intermediate adversaries generated throughout the process, retaining only the final iteration to reduce the time and space complexity.
Algorithm 1 Substitute DNN Training
Input: Training data $D=(x_{i},y_{i})_{i=1}^{n}$, simple two-layer CNN $C$, perturbation norm $\epsilon$
Initialize perturbed data $P_{0}$ with $X$
for epoch $\in\ \{1,\dots,N_{ep}\}$ do
  Train $C$ on $P_{epoch-1}$
  Update $P_{epoch}\leftarrow\{\bigl(\tilde{x}+\epsilon\cdot\text{sgn}(J_{C}(\tilde{x}),y\bigr):(\tilde{x},y)\in P_{epoch-1}\}$
end for
Return $P_{N_{ep}}$
Table 1: Description of FCNs.
Network Number of neurons in hidden layer
Net1 [100]
Net2 [60, 40]
Net3 [40, 60]
Net4 [196, 392]
Net5 [20, 20, 20, 20, 20]
4 Empirical Results
Our numerical experiments were designed to investigate the performance of our proposed correlation method across two scenarios: small neural networks and large neural networks. This approach allowed us to comprehensively evaluate the method’s efficacy and scalability.
        Refer to caption Refer to caption Refer to caption Refer to caption                     Refer to caption Refer to caption Refer to caption Refer to caption                  Refer to caption Refer to caption Refer to caption Refer to caption Before attack                     Refer to caption Refer to caption Refer to caption Refer to caption After attack
Figure 1: Test images and their predicted classes before and after attack.
Table 2: Correlation between FCNs.
(a) MNIST Net1 Net2 Net3 Net4 Net5 Net1 1.000 0.644 0.646 0.642 0.576 Net2 X 1.000 0.698 0.699 0.581 Net3 X X 1.000 0.733 0.575 Net4 X X X 1.000 0.572 Net5 X X X X 1.000 (b) Fashion-MNIST Net1 Net2 Net3 Net4 Net5 Net1 1.000 0.200 0.207 0.214 0.167 Net2 X 1.000 0.278 0.301 0.208 Net3 X X 1.000 0.289 0.204 Net4 X X X 1.000 0.221 Net5 X X X X 1.000
(c) Kuzushiji-MNIST Net1 Net2 Net3 Net4 Net5 Net1 1.000 0.564 0.577 0.590 0.523 Net2 X 1.000 0.613 0.620 0.527 Net3 X X 1.000 0.671 0.544 Net4 X X X 1.000 0.510 Net5 X X X X 1.000 (d) Afro-MNIST Net1 Net2 Net3 Net4 Net5 Net1 1.000 0.621 0.613 0.628 0.465 Net2 X 1.000 0.700 0.712 0.551 Net3 X X 1.000 0.736 0.556 Net4 X X X 1.000 0.547 Net5 X X X X 1.000
Table 3: Correlation between CNN and other networks.
Dataset Net1 Net2 Net3 Net4 Net5
MNIST 0.921 0.905 0.930 0.893 0.562
Fashion-MNIST 0.412 0.407 0.428 0.453 0.308
Kuzushiji-MNIST 0.925 0.917 0.949 0.942 0.549
Afro-MNIST 0.777 0.794 0.839 0.840 0.348
4.1 A. Small neural networks
We evaluated our method on five simple fully connected networks (FCNs) and a two-layer convolutional neural network (CNN). The number of neurons considered for each hidden layer is described in Table 1. These networks were trained on four datasets including MNIST mnist, Fashion-MNIST fashion, Kuzushiji-MNIST kuz, and Afro-MNIST afro. All these datasets are comprised of 60000 training examples and 10000 test examples. Each example consists of a $28\times 28$ image paired with a classification label from a set of 10 categories. MNIST, Fashion-MNIST, Kuzushiji-MNIST, and Afro-MNIST contain images of handwritten numerals, fashion, Hiragana characters, and African numbers respectively. The selection of datasets with similar attributes was deliberate as our focus was on understanding the interplay between networks, irrespective of data size or category.
Table 4: Accuracy of FCNs before and after attack.
Network Metric MNIST Fashion-MNIST Kuzushiji-MNIST Afro-MNIST
Net1 Accuracy before attack 96.2% 86.8% 84.4% 99.2%
Accuracy after attack 73.7% 59.5% 50.4% 45.1%
Relative degradation 23.4% 31.5% 40.3% 54.5%
Net2 Accuracy before attack 96.6% 86.1% 84.4% 99.4%
Accuracy after attack 65.7% 56.9% 47.4% 44.3%
Relative degradation 32.0% 33.9% 43.8% 55.4%
Net3 Accuracy before attack 96.0% 84.2% 83.8% 99.1%
Accuracy after attack 69.4% 53.3% 48.2% 44.5%
Relative degradation 27.7% 36.7% 42.5% 55.1%
Net4 Accuracy before attack 96.1% 86.8% 84.5% 99.2%
Accuracy after attack 67.4% 58.1% 47.2% 45.0%
Relative degradation 29.9% 33.1% 44.1% 54.6%
Net5 Accuracy before attack 92.9% 77.5% 74.3% 98.5%
Accuracy after attack 53.5% 59.5% 39.6% 40.3%
Relative degradation 42.4% 23.2% 46.7% 59.1%
CNN Accuracy before attack 98.9% 89.9% 91.7% 100.0%
Accuracy after attack 70.8% 59.5% 33.9% 27.1%
Relative degradation 28.4% 33.8% 63.0% 72.9%
Table 5: Partial correlation between ResNets.
ResNet-18 ResNet-34 ResNet-50 ResNet-101 ResNet-152
ResNet-18 1 0.661 0.288 0.312 0.133
ResNet-34 X 1 0.480 0.402 0.258
ResNet-50 X X 1 0.206 0.113
ResNet-101 X X X 1 0.052
ResNet-152 X X X X 1
Table 4 depicts the robustness of FCNs against the black-box attack detailed in Algorithm 1. The impact of the attack on some test images is illustrated in Figure 1. Results for CNN are also provided, showcasing its inferior robustness, given its role in generating the adversaries. In our specific scenario, we selected a norm value of $\epsilon=\frac{10}{255}$ and $N_{ep}=10$. The results indicate that, on average, Net5, Net2, Net3, Net4, and Net1 ranked from vulnerable to robust. Among the FCNs, Net5 exhibited the lowest overall accuracy. Notably, according to Table 2, Net5 displayed the weakest correlation with the other networks, aligning with its subpar performance and larger number of layers. Conversely, Net4 demonstrated the highest average correlation score with the other networks, potentially attributed to its larger number of neurons per layer.
In Table 3, the correlation between FCNs and CNN is investigated. Here, due to the large number of neurons involved, we had to limit the correlation analysis to only $10$ test data points. Looking at the correlation values between CNN and the different networks, Net3 and Net4 consistently demonstrates the highest correlation values across multiple datasets. This high correlation between the networks suggests that their predictive patterns and feature extraction align closely across various datasets. In addition, it’s noteworthy that Net5 exhibits the lowest correlation among the networks, which aligns with expectations due to its inferior performance and robustness.
4.2 B. Large neural networks
The proposed correlation metric was also investigated on large neural networks such as ResNet resnet, DenseNet densenet, and EfficientNet tan2019efficientnet. ResNet, short for Residual Network, is a deep neural network architecture that leverages residual blocks and skip connections to enhance training efficiency and performance. The variants of ResNet include ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 resnet. DenseNet, short for Dense Convolutional Network, is a deep neural network architecture that leverages dense connections between layers to promote feature reuse and efficient gradient flow, resulting in compact models that excel in various computer vision tasks. The architecture has several variants, namely DenseNet-121, DenseNet-161, DenseNet-169, and DenseNet-201 densenet. EfficientNet scales width, depth, and resolution proportionally to build efficient models (B0-B7) that achieve better accuracy with fewer parameters than other CNN models tan2019efficientnet. For all cases, we utilized the pretrained weights on ImageNet imagenet as our baselines. We limited our analysis to just 10 test data points sampled from the validation set of ImageNet due to the time and space complexities inherent in our methodology. Since investigating the correlation on all of the neurons in all layers was not possible, we focused our analysis solely on the initial output of the fourth layer for all ResNets, the primary output of the third transition layer for all DenseNets, and output of the third stage for all EfficientNets. A point to be highlighted is that assessing the correlation metric on the final layers often yields superior accuracy compared to other layers, given that these layers encapsulate more profound and meaningful representations. The results shown in Table 5 suggest that the most similar networks to ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 are ResNet-34, ResNet-18, ResNet-34, ResNet-34, and ResNet-34 respectively. As demonstrated in Table 6, DenseNet-161, DenseNet-121, DenseNet-201, and DenseNet-121 have the most similarities to DenseNet-121, DenseNet-161, DenseNet-169, and DenseNet-201. Similarly, Table 7 shows high correlation between adjacent scales. This pattern illustrates that networks with roughly the same number of layers often exhibit more profound partial correlations, serving as a validation of the effectiveness of our proposed correlation metric. Furthermore, deeper neural networks exhibit greater robustness against adversarial attacks. Thus, our findings further reveal that a stronger correlation with a robust model correlates with higher robustness.
Table 6: Partial correlation between DenseNets.
DenseNet-121 DenseNet-161 DenseNet-169 DenseNet-201
DenseNet-121 1 0.780 0.718 0.766
DenseNet-161 X 1 0.697 0.748
DenseNet-169 X X 1 0.725
DenseNet-201 X X X 1
Table 7: Partial correlation between EfficientNets.
EfficientNet-B0 EfficientNet-B1 EfficientNet-B2 EfficientNet-B3 EfficientNet-B4
EfficientNet-B0 1 0.820 0.811 0.816 0.819
EfficientNet-B1 X 1 0.826 0.820 0.821
EfficientNet-B2 X X 1 0.819 0.814
EfficientNet-B3 X X X 1 0.822
EfficientNet-B4 X X X X 1
5 Discussion
These experimental results indicate that our proposed correlation score can effectively demonstrate connections and relationships across diverse neural networks. While these findings are still preliminary, we assert that this score can potentially offer insights into various factors including:
  • Architectural size impact: Shallower architectures tend to exhibit higher correlation with other networks due to their simplicity and shared properties with most neural networks. A greater number of neurons per layer can contribute to a higher correlation score, as it allows for the recognition of a broader range of similar patterns with increased neural capacity.
  • Layer depth impact: Calculating partial correlations can yield more interpretable insights. Evaluation of the partial correlation metric on the earlier layers is typically less accurate since the final layers tend to capture more sophisticated and significant representations of the data.
  • Performance implications: Lower correlations may correlate with poorer model performance, indicating that neuron activations deviate from the expected patterns necessary for accurate decision-making.
  • Robustness: A strong correlation with a robust model may suggest a level of robustness within another model as well.
We further note that despite the valuable insights provided by the proposed correlation method, a principle drawback lies in the suboptimal performance in terms of time complexity for obtaining this score. Moreover, when dealing with larger models, a more efficient approach for calculating correlation with improved time complexity and precision may be necessary. Another limitation of our current methodology is the inability to pinpoint the exact reasons behind a low correlation score. However, if a high correlation is observed between two networks, where one is recognized for its accuracy and robustness, there is a strong indication that the other network is also performing well and likely to generalize. Nevertheless, this inference is not entirely precise.
6 Conclusion
We introduced a simple neuronal correlation score that compares a candidate model to a well audited reference using only forward activations on a small, unlabeled probe set. The score matches units across models with a depth aware penalty and averages the result symmetrically. Across MNIST family networks and ImageNet pretrained architectures, higher representational alignment coincides with favorable validity indicators and with robustness trends under transferable attacks. These findings support the use of representational alignment as an external, training data independent signal that complements developer controlled evaluations.